OpenStudy (anonymous):

Help please? It's my last question for the night! Glen is comparing three investment accounts offering different rates. Account A: APR of 5.80% compounding monthly Account B: APR of 5.90% compounding quarterly Account C: APR of 5.85% compounding daily Which account will give Glen at least a 6% annual yield? Account A Account B Account C Account B and Account C

OpenStudy (anonymous):

http://openstudy.com/study#/updates/5192fe22e4b0e9fd5329c180 this could help, it's the same question

OpenStudy (anonymous):

I looked at that already, couldn't figure it out :/

OpenStudy (anonymous):

Helpp :/

OpenStudy (anonymous):

ok we can do this

OpenStudy (anonymous):

alrighty keep talkin'

OpenStudy (anonymous):

for the first one , compute \[(1+\frac{.058}{4})^4-1\]

OpenStudy (anonymous):

i get this http://www.wolframalpha.com/input/?i=%281%2B.058%2F4%29^4-1 or \(0.59\) which is just under \(6\%\)

OpenStudy (anonymous):

oh damn i made a mistake, it is MONTHLY not quarterly let me start again

OpenStudy (anonymous):

I got 6% too

OpenStudy (anonymous):

..awk lol same mistake

OpenStudy (anonymous):

\[(1+\frac{.058}{12})^{12}-1\]

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=%281%2B.058%2F12%29^%2812%29-1 still just under \(6\%\) lets try the next one

OpenStudy (anonymous):

\[(1+\frac{.059}{4})^4-1\]

OpenStudy (anonymous):

I feel like it's B and C

OpenStudy (anonymous):

lets try C

OpenStudy (anonymous):

\[(1+\frac{.0585}{365})^{365}-1\]

OpenStudy (anonymous):

yeah that is the right one http://www.wolframalpha.com/input/?i=%281%2B.0585%2F365%29^%28365%29-1 it is just over \(6\%\)

OpenStudy (anonymous):

Oh! so it's both B and C for sure?

OpenStudy (anonymous):

no B is just a bit too small

OpenStudy (anonymous):

ohh so just C?

OpenStudy (anonymous):

ok i lied again let me scroll up and see if i am doing it right now i got B as a good answer

OpenStudy (anonymous):

so it has to be both right?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=%281%2B.059%2F4%29^%284%29-1 yeah B is good too

OpenStudy (anonymous):

both B and C

OpenStudy (anonymous):

yes! thank you